jack is thinking of a number. eight more than the product of 5 and the number is 93. find jack's number

Answer by ikleyn(52776)

.
Let x be the number.


5x + 8 = 93


5x = 93 - 8 

5x = 85

x = 85/5 = 17.   ANSWER